find the general solution of the equation 2x(y+z^2)p+y(2y+z^2)q=z^3. and deduce that

yz(z^2+yz-2y)=x^2 is a solution


1
Expert's answer
2021-04-15T07:25:30-0400
"2x(y+z^2)\partial z/\partial x+y(2y+z^2)\partial z/\partial y=z^2""dx\/(2x(y+z^2))=dy\/(y(2y+z^2))=dz\/z^2""ydx\/(2xy^2+2xyz^2)=xdy\/(2xy^2+xyz^2))=xydz\/(xyz^2)""(ydx-xdy-xydz)\/(2xy^2+2xyz^2-2xy^2-xyz^2-xyz^2)=(ydx-xdy-xydz)\/0""ydx-xdy-xydz=0""dx\/x-dy\/y-dz=0""lnx-lny-z=c""z=ln(x\/y)+C"

2)


"yz(z^2+2z-2y)=x^2""yz^3+2yz^2-2y^2z=x^2"

Differentiate respect to x:



"3yz^2p+4yzp-2y^2p=2x""py(3z^2+4z-2y)=2x"

Differentiate respect to y:



"z^3+3z^2yq+2z^2+4yzq-4yz-2y^2q=0""qy(3z^2+4z-2y)+z^3+2z^2-4yz=0"

So we have:



"3z^2+4z-2y=2x\/(py)""2xq\/p+z^3+2z^2-4yz=0"

Since



"p=1\/x""q=-1\/y"

then:



"-2x^2\/y+z^3+2z^2-4yz=0""-2x^2+y(z^3+2z^2-4yz)=0""yz(z^2+2z-4y)=2x^2"

So:



"(z^2+2z-4y)\/(z^2+2z-2y)=2""z^2+2z-4y=2z^2+4z-4y""z^2+2z=0"

So we get that the statement can be proved if



"z=0"

or



"z=-2"